IT IS ASKING TO SOLVE IN ELIMINATION METHOD.
5x+6Y=2
10x+12y=4
What is the solution of the system
type an ordered pair, type an integer or a fraction N if there is no solution i if there are infinate solutions

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

multiplying the 1st equation by two, gives the 2nd equation
___ this means the equations are essentially the SAME equation

so there are an infinite number of solutions

Answer by stanbon(75887)

SOLVE IN ELIMINATION METHOD.
5x+6Y=2
10x+12y=4
---------------
Multiply thru the 1st equation by 2 to get:
10x+12y = 4
----
Subtract that from the 2nd equation and you get:
0 = 0
---------------
So what is the solution?
The two equations are really the same equation.
Every point on the line 5x+6y=2 is a part of the solution.
Solution: (x, (-5/6)x + (1/3))
==================================
Cheers,
Stan H.